OpenStudy (anonymous):

what is the solution to the (3*10^2)*(5*10^5) written in scietific notation 1.5*10^8 15*10^8 1.5*10^7 15*10^7

OpenStudy (anonymous):

@Hero @helppls123 @Hannah28 @Niall @SolomonZelman

OpenStudy (anonymous):

@ranga

OpenStudy (destinymasha):

first, you need to know what the numbers with the powers represent -- 10^2 = 100. so 3 x 100 = 300. [10^2 = 10 x 10] 10^5 = 100,000 , so 5 x 100,000 = 500,000 since 10^5 = 10x10x10x10x10 NOW - 300 x 500,000 = 150000000. Which is equal to 1.5*10^8

OpenStudy (destinymasha):

Hope this helps :]